OpenStudy (anonymous):

What is the solution of n^2-49=0? -7 7 +-7 No solution

OpenStudy (anonymous):

Simplifying n2 + -49 = 0 Reorder the terms: -49 + n2 = 0 Solving -49 + n2 = 0 Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '49' to each side of the equation. -49 + 49 + n2 = 0 + 49 Combine like terms: -49 + 49 = 0 0 + n2 = 0 + 49 n2 = 0 + 49 Combine like terms: 0 + 49 = 49 n2 = 49 Simplifying n2 = 49 Take the square root of each side: n = {-7, 7}